OpenStudy (anonymous):

Solve for x if: 4x - 3(x - 2) = 4

OpenStudy (anonymous):

4x-3(x-2) = 4 distribute 3 4x-3x-6=4 simplify x-6=4 +6 to each side x=10

OpenStudy (anonymous):

I got that X = -10 ??

OpenStudy (anonymous):

Do you mind posting your work?